# thanks: These authors co-supervised the projectthanks: These authors co-supervised the project

### v1

@@ Abstract
Figure 4: Preparation of a squeezed cat state. (a) Wigner functions of the memory at early times (0, 16 ns, or 24 ns) during the deflation of |$\ket{C_{\alpha}^{+}}$, where $\alpha=2.9$. The measurements (left) are compared with numerical simulations made without free parameters (right). (b) Schematic representation of the buffer and memory states, in the non-adiabatic regime, at the beginning of deflation. An effective displacement of mode $\hat{b}$ creates an effective squeezing Hamiltonian acting on the memory during its decay to the {|$\left\{\ket{0},\ket{1}\right\}$ manifold.
@@
Interestingly, probing the deflation dynamics at short times (a few $\kappa_{b}^{-1}$) reveals a potential resource for state preparation. The deflation spontaneously prepares squeezed cat states. After preparing the state |$\ket{C_{\alpha}^{+}}$ in the memory, the drive $\epsilon_{d}(\alpha)$ is turned off while the flux remains at $\phi_{\mathrm{on}}$. The measured Wigner functions of the memory state are shown in Fig. 4a for three different times between 0 and 24 ns. As the memory releases photon pairs to the buffer, it acts as an effective drive on the buffer with an amplitude $g_{2}\hat{m}^{2}\right>=g_{2}\alpha^{2}$. This drive, which should not have an impact on the buffer when $\kappa_{b}\gg 8g_{2}\alpha$, here displaces the buffer into a state whose mean amplitude $\langle\hat{b}\rangle$ is proportional to $\alpha^{2}$. In turn an effective squeezing Hamiltonian
@@
acts on the memory, which compresses the superposition of coherent states (Fig. 4b). As the memory deflates, it reaches a maximal squeezing at a time that decreases with the initial photon number. This observation is compatible with the expected evolution of the squeezing parameter $r$ at small time in $(g_{2}|\alpha|t)^{2}$ [24]. From there, the buffer decays to vacuum, hence suppressing the effective squeezing Hamiltonian. To extract the squeezing ratio $e^{2r}$, we fit the two measured peaks in the Wigner function with Gaussian functions of identical variance. We observe squeezing ratios as high as $e^{2r}\approx 3.9~{}\mathrm{dB}$ [39], obtained after a time $t=16$ ns, starting from |$\ket{C_{\alpha}^{+}}$ with $\alpha=2.9$ (Fig. 4a). Note that during the evolution of the memory, the product of the largest and smallest variances remains close to that of the vacuum state. The predicted evolution using the bipartite buffer-memory system, without the adiabatic elimination, reproduces qualitatively the measured Wigner functions (right panels). A rounding of the Wigner function appears in the theory owing to two-photon dissipation rate, which ultimately overcomes the squeezing rate as the buffer is depleted. While bigger compression is expected starting with larger values of $\left|\alpha\right|$, an uncontrolled deformation of the initial cat state in the experiment limits the squeezing factor. We attribute this deformation and the discrepancy between simulations and experiments to the self-Kerr, the dephasing of the buffer or a detuning between $2\omega_{m}$ and $\omega_{b}$.

### v2

@@ IV Preparing squeezed cat states
Figure 4: Preparation of a squeezed cat state. (a) Pulse sequence. (b) Wigner functions of the memory at early times (0, 16 ns, or 24 ns) during the deflation of |$\ket{C_{\alpha}^{+}}$, where $\alpha=2.9$. The measurements (left) are compared with numerical simulations made without free parameters (right). (c) Schematic representation of the buffer and memory states, in the non-adiabatic regime, at the beginning of deflation. An effective displacement of mode $\hat{b}$ creates an effective squeezing Hamiltonian acting on the memory during its decay to the {|$\left\{\ket{0},\ket{1}\right\}$ manifold.
@@
Interestingly, probing the deflation dynamics at short times (a few $\kappa_{b}^{-1}$) reveals a potential resource for state preparation. The deflation spontaneously prepares squeezed cat states. After preparing the state |$\ket{C_{\alpha}^{+}}$ in the memory, the drive $\epsilon_{d}(\alpha)$ is turned off while the flux remains at $\phi_{\mathrm{on}}$ (Fig. 4a). The measured Wigner functions of the memory state are shown in Fig. 4b for three different times between 0 and 24 ns. As the memory releases photon pairs to the buffer, it acts as an effective drive on the buffer with an amplitude $g_{2}\hat{m}^{2}\right>=g_{2}\alpha^{2}$. This drive, which should not have an impact on the buffer when $\kappa_{b}\gg 8g_{2}\alpha$, here displaces the buffer into a state whose mean amplitude $\langle\hat{b}\rangle$ is proportional to $\alpha^{2}$. In turn an effective squeezing Hamiltonian
@@
acts on the memory, which compresses the superposition of coherent states (Fig. 4c). As the memory deflates, it reaches a maximal squeezing at a time that decreases with the initial photon number. This observation is compatible with the expected evolution of the squeezing parameter $r$ at small time in $(g_{2}|\alpha|t)^{2}$ (see Sec. D). From there, the buffer decays to vacuum, hence suppressing the effective squeezing Hamiltonian. To extract the squeezing ratio $e^{2r}$, we fit the two measured peaks in the Wigner function with Gaussian functions of identical variance. We observe squeezing ratios as high as $e^{2r}\approx 3.96\pm 0.07~{}\mathrm{dB}$ [38], obtained after a time $t=16$ ns, starting from |$\ket{C_{\alpha}^{+}}$ with $\alpha=2.9$ (Fig. 4b). Note that during the evolution of the memory, the product of the largest and smallest variances remains close to that of the vacuum state. The predicted evolution using the bipartite buffer-memory system, without the adiabatic elimination, reproduces qualitatively the measured Wigner functions (right panels). A rounding of the Wigner function appears in the theory owing to two-photon dissipation rate, which ultimately overcomes the squeezing rate as the buffer is depleted. While bigger compression is expected starting with larger values of $\left|\alpha\right|$, an uncontrolled deformation of the initial cat state in the experiment limits the squeezing factor. We attribute this deformation and the discrepancy between simulations and experiments to the self-Kerr, the dephasing of the buffer or a detuning between $2\omega_{m}$ and $\omega_{b}$.
V Conclusion